OpenStudy (anonymous):

Which equation will result if the substitution method is used to solve the following system of linear equations? x – y = 7 x = y + 3 A. x – y = y + 3 B. x = (7 – x) + 3 C. x – (y + 3) = 7 D. (y + 3) – y = 7

OpenStudy (anonymous):

x - y = 7 x = y+3 (y+3) - y = 7

OpenStudy (anonymous):

If x = (y+3), replace x with (y+3)